# RealSR-R1: Reinforcement Learning for Real-World Image Super-Resolution with Vision-Language Chain-of-Thought

### v2

@@ 3.3 VLCoT-GRPO
Format Reward.
To enforce the structural integrity of the multi-step reasoning process in VLCoT, we introduce a format reward that guides the model to generate outputs with predefined tags (e.g., <degradation>…</degradation>…). We use rule-based checks to ensure each step is properly wrapped in corresponding tags with correct order and completeness. A reward of $1$ is given for well-formed outputs; otherwise, 00.
Degradation Reward.
The Degradation Reward is used to supervise the model’s output and ensure the correctness of the degradation estimation. Specifically, we compare the output with the degradation ground truth obtained from the cold-start phase degradation estimation pipeline. This comparison checks whether each discrete degradation components are correctly identified and assigns scores based on the accuracy of the degradation output. The degradation reward $R_{deg}$ can be formalized as $R_{deg}=\frac{D_{correct}}{D_{total}}$, where $D_{correct}$ and $D_{total}$ represent the number of correctly estimated degradations and the total number of discrete degradation components, respectively.
Understand Reward.
In VLCoT, the coarse understanding plays a key role in identifying the main objects and layout of the image. Once the main objects are accurately recognized at this step, the subsequent middle- and fine-detail restoration steps can focus on detail enhancement and quality refinement based on the established coarse understanding and image. To encourage the model to build a clear and comprehensive global understanding early on, we design the Understand Reward $R_{und}$, which measures how well the coarse understanding text covers the main objects in the input image. Specifically, we employ Grounded-SAM [49], which supports open-vocabulary detection and segmentation, to detect main objects in the input LR images. This process extracts a set of key objects present in the scene. Subsequently, we check whether these main objects are mentioned in the coarse understanding text. If an object is explicitly referenced in the text, it is considered a hit. The Understanding Reward is defined as follows:$R_{und}=\frac{N_{cover}}{N_{total}}$, where $N_{cover}$ denotes the number of main objects correctly mentioned in the coarse understanding text, and $N_{total}$ represents the total number of main objects.
Generation Reward.
In contrast to the rule-based reward design in GRPO, the evaluation of image restoration quality in the RealSR task typically involves multiple dimensions, including perceptual visual quality, alignment, and multi-scale structural integrity. A single predefined rule is insufficient to capture these diverse aspects. To comprehensively assess the image restoration quality, we leverage a powerful prior visual expert model (e.g., Qwen2.5-VL [47]) as a visual expert to conduct a multi-faceted evaluation of the output images, considering factors such as quality, realism, and consistency. The detailed visualization of the scoring process can be found in the Appendix A.3.
@@ 4.1 Experimental Setting
Training and Testing Datasets.
For cold-start stage, we adopt the full LSDIR dataset [50] and the first 10k face images from the FFHQ dataset [51] as training dataset. Then, based on Equ. 1, we select 500 high-quality images and text pairs for training our RealSR-R1. To generate the low-quality (LR) images, we employ the degradation process of Real-ESRGAN [6] and use the same degradation settings as in SeeSR [10] to ensure a fair comparison. For testing datasets, following previous works [52, 53, 54, 13], we use three real-world datasets and one synthetic dataset for evaluation. For the real-world datasets, we test on the Dreal [55] and real [56] testset provided by Osediff [13], as well as the RealLR250 [57] dataset. Due to space limitations, the results of Dreal can be found in the Appendix A.5. Following PURE [15], we use the OST-Val [58] dataset, processed with the same degradation process in SeeSR [10], as the synthetic test set.
Table 1: Quantitative comparison with state-of-the-art real-world SR methods on both synthetic and real-world benchmarks. Best and second best performance are highlighted in red and blue, respectively.
@@
Metrics.
To comprehensively evaluate the performance of different methods, we combine full-reference and no-reference metrics. PSNR and SSIM [62] (calculated on the Y channel in the YCbCr color space) are full-reference fidelity measures, while LPIPS [63] and DISTS [64] are full-reference perceptual quality metrics. FID [65] assesses the distributional distance between the original and reconstructed images. For blind quality assessment of no-reference images, we combine NIQE [66], MANIQA [67], MUSIQ [68], CLIPIQA [69], and TOPIQ [70] to evaluate the naturalness and structural integrity of the images.
Implementation Details.
In the cold-start phase, we use the AdamW optimizer with an initial learning rate of $2\times 10^{-5}$, and employ a cosine annealing strategy for learning rate decay. The model is trained for 2 epochs on the entire dataset with batchsize=16, including a warmup phase spanning the first 0.01 epochs to stabilize early-step optimization. In the VLCoT-GRPO phase, the AdamW optimizer is again used with an initial learning rate of $1\times 10^{-5}$, a batch size of 8, and a constraint on the maximum generated token length of $6k$. For the same LR image, RealSR-R1 samples three responses and provides corresponding rewards. The model was trained for 2 epochs.
@@ 4.2 Comparison with State-of-the-Arts
Quantitative Comparisons.
The quantitative comparison is shown in Tab. 1. On synthetic datasets, RealSR-R1 achieves the best or second-best results across seven metrics, reflecting the capability to generate high-quality and realistic images. On real-world benchmarks, our method performs strongly across most no-reference metrics and shows comparable performance to diffusion-based methods on reference-based metrics. However, PSNR and SSIM still have a gap with diffusion-based methods. This is due to the trade-off between realism and fidelity, as RealSR-R1 generates more textures and details, which may reduce fidelity metrics. Recent studies [71, 9, 57] have shown that the restored images exhibit higher quality in terms of human perception, but perform poorly on certain reference-based metrics (e.g., PSNR and SSIM). Therefore, it is necessary to reconsider the reference values of existing metrics and propose more effective methods to evaluate modern image restoration methods.
@@
Qualitative Comparisons.
In Fig. 3, we present a visual comparison of different datasets. As shown in the first case, even when faced with severely degraded input, RealSR-R1 generates realistic wall texture details, while other methods produce overly smooth or incorrect textures. In the second example, SeeSR and OSEDiff produce blurred structures in the background buildings, while PURE introduces excessive pseudo-textures. In contrast, our method most accurately restores the image details. On the RealLR250 dataset, RealSR-R1 also demonstrates an advantage in visual quality, generating more realistic face details while maintaining better overall image fidelity. Additional visual comparison results and the complete CoT output are provided in the Appendix A.4.
User Study.
To further demonstrate that the images generated by our method align more closely with human perception, we conducted a user study with 20 participants, evaluating both synthetic and real-world data. Specifically, we randomly selected 50 real-world LR images from RealSR, DrealSR, and RealLR250, and chose 25 degraded images from the synthetic dataset OST-Val. We then compared the enhancement results of RealSR-R1 with six methods (PURE, SeeSR, StableSR, PASD, DiffBIR, and OSEDiff). Participants were asked to vote on the best restoration results based on visual quality, naturalness, and accuracy. As shown in Fig. 4, our RealSR-R1 achieved better preferences across different datasets than other methods, with a selection rate of 30.7% and 43.6%, respectively. This result demonstrates its advantage in generating high-quality images.
@@
Understanding Capabilities of Different Models
We further compared the understanding capabilities of GAN-based, Diffusion-based, and MLLM-based methods for real image super-resolution. As shown in Fig. 5, GAN-based methods lack an understanding of image content, while diffusion-based methods utilize external language models (e.g., DAPE [10]) to extract simple image descriptions as prompts. MLLMs possess more refined understanding capabilities, but directly applying this fine-grained understanding to severely degraded images leads to incorrect interpretations of details. For example, diffusion and MLLM mistakenly interpret the image as a bear and a rabbit, respectively. In contrast, only RealSR-R1 correctly recognizes the image as a raccoon and generates the most accurate and detailed image description, which is crucial for restoring image details.
@@ 4.3 Ablation Study
Table 2: Ablation of the VLCot on OST-Val.
@@
Table 3: Ablation on reward functions.
@@
Ablation of the key components of RealSR-R1.
To validate the effectiveness of RealSR-R1, we systematically conducted ablation experiments on OST-Val. We progressively removed core components from RealSR-R1. The results, as shown in Tab 3, removing the VLCoT GRPO leads to significant reductions in all metrics, demonstrating the substantial potential of R1 in real-world super-resolution. Secondly, the degradation and understanding output affect different metrics: removing the degradation output and removing the understanding output significantly reduce the reference-based metrics SSIM, LPIPS, and the no-reference metric MANIQA, respectively. Furthermore, removing multi-step outputs and reducing the length of CoT performs poorly on most metrics. These observations validate the effectiveness of each core component of the proposed method.
Ablation of VLCoT-GRPO and reward functions.
We further performed a step-by-step removal of the proposed reward functions to ablate their effectiveness. In Tab. 3, we find that different rewards affect different metrics. For example, the generation reward significantly improves CLIPIQA and TOPIQ, while the degradation reward tends to improve SSIM and LPIPS. Additionally, removing the understanding reward impacts image restoration quality across all metrics, indicating that semantic understanding plays a crucial role in driving the generation of real-world details.
5 Conclusion and Limitation
Appendix A Appendix
A.1 More Ablations
The ablation of understanding reward.
We conduct ablation studies on tag extraction strategies within the understanding reward to evaluate the impact of different tag extraction methods on model performance. As shown in Tab 4, RAM [72] performs poorly on both perceptual and generative metrics, which may be attributed to the fact that the extracted tags often contain incorrect or misleading information. In contrast, the tags extracted by Ground-SAM achieve the best performance across all four metrics. Therefore, we adopt Ground-SAM as the tag extraction method for the understanding reward in RealSR-R1.
Table 4: Ablation of tag extraction strategies within the Understanding reward.
@@ A.1 More Ablations
The ablation of generation reward.
We also conduct ablation studies on the visual expert used in the generation reward. As shown in Tab. 5, employing either perceptual metrics or generative metrics as the visual expert significantly enhances the corresponding perceptual or generative quality—while simultaneously degrading performance on the other. In contrast, Qwen2.5-VL achieves the best trade-off between these two objectives. Therefore, we adopt Qwen2.5-VL as the visual expert in RealSR-R1 to score generated images. It is also worth noting that using hand-crafted metrics involves evaluating each score independently, whereas Qwen2.5-VL can jointly assess the perceptual and generative quality of a given image set, leading to more holistic and context-aware scoring. Visualization of the scoring process is provided in SecA.3.
Table 5: Ablation of the generation reward. The perceptual metric is computed as a weighted combination of PSNR, SSIM, LPIPS, and DISTS scores, while the generative quality metric is defined as a weighted aggregation of CLIPIQA, MANIQA, MUSIQ, and TOPIQ scores.
@@
Perceptual metrics 0.4298 0.3672 0.6578 0.6849 0.7633
Generate metrics 0.3595 0.4781 0.6869 0.7106 0.8168
Qwen2.5-VL 0.3963 0.4117 0.6858 0.7127 0.8086
A.2 The Detailed Prompts Used in the Dialogue.
we leverage Qwen2.5-VL [47] to generate ground-truth supervision for textual understanding at coarse, middle, and fine levels—required during the cold-start stage. As illustrated in Fig. 6, we employ a three-round dialog process, where progressively detailed image descriptions are generated through prompt-based interaction. Each round is guided by specifically designed prompts to elicit increasingly fine-grained textual understanding of the visual content.
A.3 Scoring Process for Generating Rewards.
Fig. 7 illustrates the scoring process of generated images using a powerful prior visual expert model (e.g., Qwen2.5-VL [47]). The expert conducts a multi-faceted evaluation of each output image, taking into account factors such as quality, realism, and consistency. Alongside the individual scores, the expert provides interpretability through rationale for each assessment. Finally, the scores across all dimensions are aggregated via a weighted sum to produce an overall image quality score.
A.4 Visualization of the Complete VLCoT Output.
We further visualize the full image and text output process of VLCoT. As shown in Fig. 8, RealSR-R1 progressively generates multi-level textual understanding and multi-resolution image outputs, which serve as intermediate "gray" reasoning steps. This step-by-step process ultimately leads to an accurate reconstruction of the target image.
A.5 More Extensive Quantitative Results.
We further report comparative results between RealSR-R1 and state-of-the-art RealSR models on the DReal dataset. As shown in Tab 6, RealSR-R1 achieves superior performance across multiple generative metrics (e.g., NIQE [66], MANIQA [67], MUSIQ [68], CLIPIQA [69], and TOPIQ [70]), demonstrating its strong ability to generate more realistic images. It is worth noting that although VARSR achieves competitive performance in image generation, its training relies on proprietary data. In contrast, RealSR-R1 surpasses VARSR in generation metrics while being trained solely on publicly available datasets.
Table 6: Quantitative comparison with state-of-the-art real-world SR methods on both synthetic and real-world benchmarks. Best and second best performance are highlighted in red and blue, respectively.
A.6 More Visualization Results.
We add more visualization results of our RealSR-R1, making comparisons with other SOTA methods on different datasets. As shown in Fig.9 and Fig.10, RealSR-R1 achieves superior performance in scenes involving human faces, animal fur, architectural structures, and plant textures, yielding results that are more consistent with human perceptual preferences.
Figure 8: Visual example of VLCoT’s complete image and text output process.
Figure 9: Qualitative comparisons with different SOTA methods.
Figure 10: Qualitative comparisons with different SOTA methods.

### v3

@@ 3.3 VLCoT-GRPO
Reinforcement learning has been proven effective in enhancing reasoning capability in large language models, as demonstrated by the GRPO [15]. To enable cross-modal collaborative reasoning in RealSR, we propose VLCoT-GRPO. This reinforcement learning framework optimizes visual and textual reasoning processes within the VLCoT. Compared with pure text-based reasoning tasks, multi-modal CoT reinforcement presents unique challenges. Due to the inherent limitation of current training, most MLLMs cannot natively perform alternating image-text generation. The model must dynamically determine when to reason for semantic understanding or image restoration. Traditional supervised fine-tuning rely on labeled prompts specifying which modality to generate, whereas VLCoT-GRPO introduces implicit task coordination via reinforcement learning.
For a given query $q$, VLCoT-GRPO performs iterative vision–language reasoning over $T$ steps. At each step $t$, the model alternates between textual reasoning and visual reconstruction, producing an interleaved output sequence $o_{i}=\{l_{i},v_{i}\}$, where $l_{i}$ denotes the $i^{th}$ languag token and $v_{i}$ denotes the corresponding vision token. Each reasoning step refines both semantic understanding and visual reconstruction conditioned on all previous steps.
Following the GRPO framework, the policy ratio between the current and old policies at step $t$ is defined as:
$\begin{cases}\dfrac{\pi_{\theta}(l_{l,j}\mid q,o_{<t},l_{t,<j})}{\pi_{\theta_{\text{old}}}(l_{l,j}\mid q,o_{<t},l_{t,<j})},&0\leq j\leq|l_{t}|,\\[6.0pt] \dfrac{\pi_{\theta}(v_{l,j-|l_{t}|}\mid q,o_{<t},l_{t},v_{t,<j-|l_{t}|})}{\pi_{\theta_{\text{old}}}(v_{l,j-|l_{t}|}\mid q,o_{<t},l_{t},v_{t,<j-|l_{t}|})},&|l_{t}|<j\leq|l_{t}|+M_{t},\end{cases}$ (5)
where $M_{t}$ represents the number of visual tokens generated in step $t$.
3.4 Reward Functions
@@
Format Reward. To enforce the structural integrity of the multi-step reasoning process in VLCoT, we introduce a format reward that guides the model to generate outputs with predefined tags (e.g., <degradation>…</degradation>…). We use rule-based checks to ensure each step is properly wrapped in corresponding tags with correct order and completeness. A reward of $1$ is given for well-formed outputs; otherwise, 0.
Degradation Reward. The Degradation Reward is used to supervise the model’s output and ensure the correctness of the degradation estimation. Specifically, we compare the output with the degradation ground truth obtained from the cold-start phase degradation estimation pipeline. This comparison checks whether each discrete degradation components are correctly identified and assigns scores based on the accuracy of the degradation output. The degradation reward $R_{deg}$ can be formalized as $R_{deg}=\frac{D_{correct}}{D_{total}}$, where $D_{correct}$ and $D_{total}$ represent the number of correctly estimated degradations and the total number of discrete degradation components, respectively.
[…]
Generation Reward. In contrast to the rule-based reward design in GRPO, the evaluation of image restoration quality in the RealSR task typically involves multiple dimensions, including perceptual visual quality, alignment, and multi-scale structural integrity. A single predefined rule is insufficient to capture these diverse aspects. To comprehensively assess the image restoration quality, we leverage a powerful prior visual expert model (e.g., Qwen2.5-VL [2]) as a visual expert to conduct a multi-faceted evaluation of the output images, considering factors such as quality, realism, and consistency. Fig. 4 illustrates the scoring process of generated images. Alongside the individual scores, the expert provides interpretability through a rationale for each assessment. Finally, the scores across all dimensions are aggregated via a weighted sum to produce an overall image quality score. It is worth mentioning that experts judge the quality of a group of multiple restored images simultaneously. This is because we encourage the model to learn relative preferences instead of memorizing fixed quality values. This comparative learning paradigm helps RealSR-R1 develop its own internal ”value system” for generation quality, rather than simply overfitting to the biases of the reward model.
@@ 4.1 Experimental Setting
Training and Testing Datasets. For cold-start stage, we adopt the full LSDIR dataset [28] and the first 10k face images from the FFHQ dataset [23] as training dataset. Then, based on Equ. 1, we select 500 high-quality images and text pairs for training our RealSR-R1. To generate the low-quality (LR) images, we employ the degradation process of Real-ESRGAN [55] and use the same degradation settings as in SeeSR [63] to ensure a fair comparison. For testing datasets, following previous works [54, 14, 44, 62], we use three real-world datasets and one synthetic dataset for evaluation. For the real-world datasets, we test on the Dreal [61] and real [3] testset provided by Osediff [62], as well as the RealLR250 [1] dataset. Following PURE [59], we use the OST-Val [56] dataset, processed with the same degradation process in SeeSR [63], as the synthetic test set.
TABLE I: Quantitative comparison with state-of-the-art real-world SR methods on both synthetic and real-world benchmarks. Best and second best performance are highlighted in red and blue, respectively.
@@
Evaluation metrics. To comprehensively evaluate the performance of different methods, we combine full-reference and no-reference metrics. PSNR and SSIM [58] (calculated on the Y channel in the YCbCr color space) are full-reference fidelity measures, while LPIPS [76] and DISTS [12] are full-reference perceptual quality metrics. FID [18] assesses the distributional distance between the original and reconstructed images. For blind quality assessment of no-reference images, we combine NIQE [73], MANIQA [65], MUSIQ [24], CLIPIQA [53], and TOPIQ [6] to evaluate the naturalness and structural integrity of the images.
Implementation Details. In the cold-start phase, we use the AdamW optimizer with an initial learning rate of $2\times 10^{-5}$, and employ a cosine annealing strategy for learning rate decay. The model is trained for 2 epochs on the entire dataset with batchsize=16, including a warmup phase spanning the first 0.01 epochs to stabilize early-step optimization. In the VLCoT-GRPO phase, the AdamW optimizer is again used with an initial learning rate of $1\times 10^{-5}$, a batch size of 8, and a constraint on the maximum generated token length of $6k$. For the same LR image, RealSR-R1 samples three responses and provides corresponding rewards. The model was trained for 2 epochs.
@@ 4.2 Comparison with State-of-the-Arts
Quantitative Comparisons. The quantitative comparison is shown in Tab. I. On synthetic datasets, RealSR-R1 achieves the best or second-best results across seven metrics, reflecting the capability to generate high-quality and realistic images. On real-world benchmarks, our method performs strongly across most no-reference metrics (e.g., NIQE [73], MANIQA [65], MUSIQ [24], CLIPIQA [53], and TOPIQ [6]), demonstrating its strong ability to generate more realistic images. It is worth noting that although VARSR [44] achieves competitive performance in image generation, its training relies on proprietary data. In contrast, RealSR-R1 surpasses VARSR in generation metrics while being trained solely on publicly available datasets. However, PSNR and SSIM still have a gap with diffusion-based methods. This is due to the trade-off between realism and fidelity, as RealSR-R1 generates more textures and details, which may reduce fidelity metrics. Similarly, recent state-of-the-art methods such as PURE [59] and VARSR also struggle to simultaneously achieve the best PSNR and SSIM while maintaining strong perceptual quality. In contrast, RealSR-R1 provides greater controllability. By adjusting the orientation of the generation reward within VLCoT-GRPO, the model can be guided toward fidelity-oriented reconstruction. A more detailed analysis of this trade-off and the corresponding ablation experiments are presented in Tab. VII in Sec. 4.3.
Qualitative Comparisons. In Fig. 5, we present a visual comparison of different datasets. As shown in the first case, even when faced with severely degraded input, RealSR-R1 generates realistic wall texture details, while other methods produce overly smooth or incorrect textures. In the second example, SeeSR and OSEDiff produce blurred structures in the background buildings, while PURE introduces excessive pseudo-textures. In contrast, our method most accurately restores the image details. On the RealLR250 dataset, RealSR-R1 also demonstrates an advantage in visual quality, generating more realistic face details while maintaining better overall image fidelity.
User Study. To further demonstrate that the images generated by our method align more closely with human perception, we conducted a user study with 20 participants, evaluating both synthetic and real-world data. Specifically, we randomly selected 50 real-world LR images from RealSR, DrealSR, and RealLR250, and chose 25 degraded images from the synthetic dataset OST-Val. We then compared the enhancement results of RealSR-R1 with six methods (PURE, SeeSR, VARSR, PASD, DiffBIR, and OSEDiff). Participants were asked to vote on the best restoration results based on visual quality, naturalness, and accuracy. Importantly, to ensure reflects real-world deployment conditions, we did not provide the ground-truth for participants. This no-reference setup aligns with practical scenarios where end users typically have access only to the degraded input and model outputs, not the ground-truth. We believe this design better captures human perceptual preference and the practical utility of each method. As shown in Fig. 6, our RealSR-R1 achieved better preferences across different datasets than other methods, with a selection rate of 30.7% and 43.6%, respectively. This result demonstrates its advantage in generating high-quality images.
@@
Understanding Capabilities of Different Models We further compared the understanding capabilities of GAN-based, Diffusion-based, and MLLM-based methods for real image super-resolution. As shown in Fig. 7, GAN-based methods lack an understanding of image content, while diffusion-based methods utilize external language models (e.g., DAPE [63]) to extract simple image descriptions as prompts. MLLMs possess more refined understanding capabilities, but directly applying this fine-grained understanding to severely degraded images leads to incorrect interpretations of details. For example, diffusion and MLLM mistakenly interpret the image as a bear and a rabbit, respectively. In contrast, only RealSR-R1 correctly recognizes the image as a raccoon and generates the most accurate and detailed image description, which is crucial for restoring image details.
Model Complexity Comparison with PURE We have conducted a complexity and efficiency comparison between our RealSR-R1 model and PURE [59], a representative autoregressive multimodal generation method. All evaluations were performed using $512\times 512$ input images on an NVIDIA A800 GPU under consistent conditions.
TABLE II: model complexity comparisons with the output size of $512\times 512$.
Method Time(s) PSNR↑ SSIM↑ LPIPS↓ DISTS↓ FID↓
PURE 316 17.98 0.3816 0.4664 0.2317 70.74
RealSR-R1 372 18.29 0.3963 0.4117 0.2126 62.82
Param(M) NIQE↓ MUSIQ↑ MANIQA↑ CLIPIQA↑ TOPIQ↑
PURE 7080 4.9768 70.5592 0.6547 0.7038 0.7775
RealSR-R1 7080 4.1504 71.4819 0.6858 0.7127 0.8086
As shown in Tab. II, RealSR-R1 has the same number of parameters as PURE. The inference time of our model is slightly longer (+56s), primarily due to the inclusion of coarse and middle-detail stages. However, this design substantially enhances the model’s visual understanding and generation capabilities with the help of R1, leading to notable improvements across all ten evaluation metrics, which is very difficult, so the slight sacrifice in speed is acceptable. It is also worth highlighting that, to the best of our knowledge, our work is the first to successfully adapt an R1-style architecture to a low-level vision task and empirically demonstrate its effectiveness. We believe this provides a valuable foundation for future exploration of autoregressive reasoning in image restoration.
Extension to Image Deraining Task To further evaluate the generalizability of RealSR-R1 to other image restoration tasks, we conducted an experiment on image draining. Specifically, we generated a deraining dataset by adding synthetic raindrops to clean training images. We then fine-tuned RealSR-R1 for only 10K iterations (including cold-start and VLCoT-GRPO) on this task. Despite the limited training, the model still shows remarkable performance, as shown in Tab. III.
TABLE III: Result on image deraining task.
Rian100H
[…]
MPRNet [71] 30.41 0.158 69.5733 0.6788 0.6534 0.6676
IR-SDE [36] 31.65 0.047 70.6452 0.6811 0.7782 0.7085
RealSR-R1 33.59 0.419 72.3578 0.7182 0.8196 0.7741
These results demonstrate that VLCoT’s step-wise visual reasoning and the VLCoT-GRPO training strategy can benefit other image restoration tasks.
@@ 4.3 Ablation Study
TABLE IV: Ablation of the VLCot on OST-Val.
@@
TABLE V: Ablation on reward functions.
@@
Ablation of the key components of RealSR-R1. To validate the effectiveness of RealSR-R1, we systematically conducted ablation experiments on OST-Val. We progressively removed core components from RealSR-R1. The results, as shown in Tab V, removing the VLCoT GRPO leads to significant reductions in all metrics, demonstrating the substantial potential of R1 in real-world super-resolution. Secondly, the degradation and understanding output affect different metrics: removing the degradation output and removing the understanding output significantly reduce the reference-based metrics SSIM, LPIPS, and the no-reference metric MANIQA, respectively. Furthermore, removing multi-step outputs and reducing the length of CoT performs poorly on most metrics. These observations validate the effectiveness of each core component of the proposed method.
Ablation of VLCoT-GRPO and reward functions. We further performed a step-by-step removal of the proposed reward functions to ablate their effectiveness. In Tab. V, we find that different rewards affect different metrics. For example, the generation reward significantly improves CLIPIQA and TOPIQ, while the degradation reward tends to improve SSIM and LPIPS. Additionally, removing the understanding reward impacts image restoration quality across all metrics, indicating that semantic understanding plays a crucial role in driving the generation of real-world details.
The ablation of understanding reward. We conduct ablation studies on tag extraction strategies within the understanding reward to evaluate the impact of different tag extraction methods on model performance. As shown in Tab VI, RAM [78] performs poorly on both perceptual and generative metrics, which may be attributed to the fact that the extracted tags often contain incorrect or misleading information. In contrast, the tags extracted by Ground-SAM achieve the best performance across all four metrics. Therefore, we adopt Ground-SAM as the tag extraction method for the understanding reward in RealSR-R1.
TABLE VI: Ablation of tag extraction strategies within the Understanding reward.
@@
The ablation of generation reward. We further conduct ablation studies on the visual expert used in the generation reward to verify the controllability of RealSR-R1. As shown in Tab. VII, using perceptual metrics as the reward leads the model to prioritize reconstruction fidelity—achieving the highest PSNR and SSIM. In contrast, employing generative quality metrics drives the model toward visually appealing and realistic details, obtaining the best MANIQA and TOPIQ scores. These results demonstrate that RealSR-R1 possesses controllable generation bias. By adjusting the reward orientation within the VLCoT-GRPO framework, the model can flexibly balance between reconstruction fidelity and perceptual realism, thereby achieving restoration outcomes that better align with different application requirements.
Notably, when using Qwen2.5-VL as the visual expert, the model achieves a balanced performance across all indicators—retaining relatively high PSNR/SSIM while improving perceptual metrics (e.g., CLIPIQA 0.7127, MANIQA 0.6858)—representing the most effective trade-off. Moreover, unlike hand-crafted metrics that evaluate each aspect independently, Qwen2.5-VL jointly assesses perceptual and generative quality for the entire output set, providing more holistic and context-aware feedback for reinforcement learning optimization.
TABLE VII: Ablation of the generation reward. The perceptual metric is computed as a weighted combination of PSNR, SSIM, LPIPS, and DISTS scores, while the generative quality metric is defined as a weighted aggregation of CLIPIQA, MANIQA, MUSIQ, and TOPIQ scores.
@@
Perceptual metrics 19.89 0.4298 0.3672 0.6578 0.6849 0.7633
Generate metrics 17.73 0.3595 0.4781 0.6869 0.7106 0.8168
Qwen2.5-VL 18.29 0.3963 0.4117 0.6858 0.7127 0.8086
The ablation of different stage design. To further investigate the impact of our three-stage design, we conducted a comparison between three-step and two-step variants (removing the middle-detail stage). Results are shown in Tab. VIII.
TABLE VIII: Ablation of different step of RealSR-R1.
Method SSIM↑ LPIPS↓ MANIQA↑ CLIPIQA↑ TOPIQ↑
Two-step 0.3924 0.4248 0.6781 0.7106 0.7913
Three-step 0.3963 0.4117 0.6858 0.7127 0.8086
These results demonstrate that the three-step reasoning process consistently provides better performance across all metrics. We hypothesize that this is because the two-step pipeline introduces a larger semantic gap between stages, making it harder for the model to progressively refine the image in a controllable manner.
Figure 8: Comparison of detection accuracy of clean and degraded images using Grounded-SAM and Robust SAM.
Performance of Grounded-SAM on low-quality images. To evaluate whether the performance of Grounded-SAM itself will degrade on low-quality images, we randomly selected 50 images from the RealSR and OST-Val test datasets and manually annotated their main objects. We then conducted a quantitative analysis of Grounded-SAM’s detection accuracy on the test dataset RealLQ250. As shown in the Fig. 8, while Grounded-SAM does exhibit a drop in accuracy under degraded conditions, it still maintains a detection accuracy exceeding 80%. To address this limitation further, we evaluated the performance of RobustSAM [8]—a recent variant designed to enhance SAM’s robustness on degraded images. Preliminary results show that RobustSAM achieves a +6.4% improvement in accuracy over Grounded-SAM. We are actively exploring the integration of more resilient detectors to mitigate object recognition failures during training, and in the future will add confidence thresholds in reward formulation to minimize the risk of false penalties.
5 Conclusion and Limitation
Acknowledgment
This work is supported by the National Natural Science Foundation of China (NO. 62572193, 62102151), Shanghai Science and Technology Commission (22511104600), the Open Research Fund of the Key Laboratory of Advanced Theory and Application in Statistics and Data Science, Ministry of Education, and the Fundamental Research Funds for the Central Universities.